Part Specification/Six Sigma

  • A design team was charged with a Cpk = 1.3 and Nominal of Peak Pulse Current of 10μ They measured 30 parts from one of the vendors and found the average to be 10μA, with a standard deviation of 0.2. What should the part specification be set to?

Specification = 10μA +/- __?_____
Answer #1

Process capacity index is nothing but the statistical measure of process capacity to identify whether output is produced within customer's specification limits. The standard Cpk level is 2.0 for the six sigma process.

Process capacity is the number of times the process fits into the tolerance width. So, higher the value of Cp, the better the process.

CPK = 13 Mean = 10 MA o =0.2 USL 10 +ne LSL = 10-C %3D CPK = min USL - mean LSL mean 6. 30 10-(10-10)) 3X0.2 1 3 min to +C-1